OpenStudy (anonymous):

Assume that 25 percent of voters are Republicans, 45 percent are Democrats, and 30 percent are Independents. A poll taker asks 10 people at random for their party affiliation. Assume further that, because the city is so large, this situation can be approximated by a Bernoulli process. What is the probability that at least 8 of those polled are Democrat voters________________?

OpenStudy (anonymous):

at least 8 means you have work to do you have to compute \[P(x=8), P(x=9), P(x=10)\] and add there is no way around it

OpenStudy (anonymous):

put p = .45, 1- p = .55 and use the formula \[P(x=8)=\dbinom{10}{8}(.45)^8(.55)^2=45(.45)^8(.55)^2\] \[P(x=9)=10(.45)^9(.55)\] \[P(x=10)=(.45)^{10}\] compute and add

OpenStudy (anonymous):

10 and 8 in the first one is a combination right?

OpenStudy (anonymous):

yeah it is 10 choose 8 which is 45

OpenStudy (anonymous):

ok thank you again

OpenStudy (anonymous):

When I compute the second part of the equation I get 4.161743 and so on

OpenStudy (anonymous):

am I not computing it right